OpenStudy (anonymous):

A triangle has sides a b c and angles A B C. If angle A is 81, side a is 11 and side b is 12, what is angle B?

OpenStudy (anonymous):

use cosine law.

Parth (parthkohli):



Parth (parthkohli):

Do what shubham says: cosine rule.

OpenStudy (anonymous):

i.e. \[b^{2}=a^{2}+c^{2}-2ac cosB\]

OpenStudy (anonymous):

But I don't know what side c is?

OpenStudy (anonymous):

So I have two unknowns

OpenStudy (anonymous):

first find c.

OpenStudy (anonymous):

using cosine law, \[c^{2}=a^{2}+b^{2}-2ab cosC\]

OpenStudy (anonymous):

Sorry, I don't know angle C either...should I use sine rule instead?

OpenStudy (anonymous):

oh sorry, you can use sine law,

OpenStudy (anonymous):

\[\frac{sinA}{a}=\frac{sinB}{b}=\frac{}{}\]

OpenStudy (anonymous):

So...sin A/a = sin B/b, which would be sin 81/11 = sin B/12

OpenStudy (anonymous):

So (12 sin 81)/11 = sin B. sin B = 1.077

OpenStudy (anonymous):

But my calculator gives an error message when I try to find sin -1 of 1.077 - shouldn't values of sin be less than 1?

OpenStudy (anonymous):

Have I gone wrong somewhere?

OpenStudy (anonymous):

no , i am also getting 1.077 there would be problem with given data

OpenStudy (anonymous):

Oh...I just double checked the numbers on the question, and I didn't copy them wrongly...the question asks if there are two possible values for angle B, so perhaps it's a trick question and this one has no possible solutions....hmm, not sure. Thanks for your help anyway.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Hi @Kashan, how did you get sin B = 12/11? Did you use sine rule?